# T2M-HiFiGPT: Generating High Quality Human Motion from Textual Descriptions with Residual Discrete Representations

### 17 Dec 2023 v1

@@ Abstract
In this study, we introduce T2M-HiFiGPT, a novel conditional generative framework for synthesizing human motion from textual descriptions. This framework is underpinned by a Residual Vector Quantized Variational AutoEncoder (RVQ-VAE) and a double-tier Generative Pretrained Transformer (GPT) architecture. We demonstrate that our CNN-based RVQ-VAE is capable of producing highly accurate 2D temporal-residual discrete motion representations. Our proposed double-tier GPT structure comprises a temporal GPT and a residual GPT. The temporal GPT efficiently condenses information from previous frames and textual descriptions into a 1D context vector. This vector then serves as a context prompt for the residual GPT, which generates the final residual discrete indices. These indices are subsequently transformed back into motion data by the RVQ-VAE decoder. To mitigate the exposure bias issue, we employ straightforward code corruption techniques for RVQ and a conditional dropout strategy, resulting in enhanced synthesis performance. Remarkably, T2M-HiFiGPT not only simplifies the generative process but also surpasses existing methods in both performance and parameter efficacy, including the latest diffusion-based and GPT-based models. On the HumanML3D [25] and KIT-ML [58] datasets, our framework achieves exceptional results across nearly all primary metrics. We further validate the efficacy of our framework through comprehensive ablation studies on the HumanML3D [25] dataset, examining the contribution of each component. Our findings reveal that RVQ-VAE is more adept at capturing precise 3D human motion with comparable computational demand compared to its VQ-VAE counterparts. As a result, T2M-HiFiGPT enables the generation of human motion with significantly increased accuracy, outperforming recent state-of-the-art approaches such as T2M-GPT [79] and Att-T2M [81].
@@ 1 Introduction
Action generation based on natural language is challenging, mainly because (1) human motion in the real physical world is high-dimensional data with irregular spatial structure, highly dynamic temporal characteristics and many fine details. Direct text-driven synthesis in high-dimensional posture space (such as [45, 71, 69] etc.) is challenging, and slight changes to high-dimensional vectors can lead to visible defects in the generated motion. (2) Actions and text come from different modalities. The model needs to learn the complex mapping relationship from language space to action space across modalities. For this, many works propose using autoencoders [3, 24, 68, 77] and VAE [81, 79, 56, 57, 15] to learn the joint embedding of language and action. MotionClip [68] aligns the action space with the CLIP [59] space. ACTOR [56] and TEMOES [57] respectively proposed Transformer-based VAEs for action-to-action and text-to-action. These works show promising performance under simple descriptions and are limited to generating high-quality actions when the text description becomes long and complex. Guo et al. [25] and TM2T [26] aim to generate action sequences with more challenging text descriptions. However, both methods are not direct, involve three stages of text-to-action generation, and sometimes fail to generate high-quality actions consistent with the text. Recently, diffusion-based models [33] have shown impressive results in image generation [61], then introduced into action generation by MDM [69], MotionDiffuse [80] and MotionLDM [15], and dominate the text-to-action generation task. However, we find that the performance and inference speed improvement of diffusion-based methods [69, 80, 15] may not be as significant as compared to classical methods (such as VQVAE [72]). Recent T2M-GPT [79] and Att-T2M [81] work, both use the VQVAE [72] model as the motion discrete representation; however, we find that real motion capture data contain very rich details, and simply using a single layer of discrete representation is not enough to fully and compactly express these motion details. Simply increasing the number of codebooks will lead to more redundancy, and the motion representation is not compact enough, which will also increase the training difficulty of the generation model due to large codebook sizes and data scarcity. In this work, unlike T2M-GPT [79] and Att-T2M [81], we are inspired by the latest progress in learning discrete representations for generation [5, 19, 20, 22, 49, 61, 72, 75, 78, 18, 74], and studied a simple framework based on Residual Vector Quantization Variational Autoencoder (RVQ-VAE) [78] and Valle-like [74, 40] structure for the compact and highly accurate discrete motion representation and text-to-action generation task. Specifically, we present a two-stage approach for synthesizing actions from textual descriptions. Initially, we employ a standard 1D dilated convolutional network to transform the action sequence into multi-layer residual discrete code indices. Subsequently, drawing inspiration from the speech generation model Valle [74], we introduce an innovative dual-tier GPT model designed to generate motion residual code sequences from pre-trained CLIP [59] text embeddings. Our experimental analysis reveals that the RVQ-VAE model can effectively decrease the downsampling factor of motion data (from 4 to 8), while simultaneously reducing the computational time required for inference. This results in superior reconstruction outcomes on the test set. Moreover, as the sequence length for modeling is halved, the inference duration for the second stage is also proportionately diminished. Despite its straightforward nature, our method is capable of producing high-fidelity motion sequences that are intricately aligned with complex textual descriptions. In practical evaluations, our approach outperformed the GPT-based methods T2M-GPT [79] and Att-T2M [81] on the widely recognized datasets HumanML3D [25] and KIT-ML [58]. Notably, on the extensive HumanML3D [25] dataset, our method achieved the highest accuracy, even exceeding the real data in terms of the consistency between text and generated motion, as measured by R-Precision and matching score. Through comprehensive experimentation, we have delved into this domain, and we hope that the insights and conclusions drawn from our work can contribute to future advancements in the field.
@@
    Our findings indicate that the RVQ-VAE model with shared codebook, despite having comparable computational and parameter requirements to the VQ-VAE, exhibits superior performance in reconstructing 3D human motion data, thereby establishing itself as a more effective method for the discrete representation of 3D human motion data.
@@ 2 Related Work
Text-based human action generation. Text-based human action generation is a burgeoning field focused on creating 3D human actions from textual descriptions. Early work in this area includes Text2Action [2], which utilized an RNN-based model to produce actions conditioned on brief texts. Language2Pose [3] took this further by employing a curriculum learning strategy to learn a joint embedding space of text and pose, enabling the generation of action sequences from text embeddings. Subsequent research by Ghosh et al. [24] improved upon this by learning separate manifold representations for upper and lower body movements, yielding better results than Language2Pose [3]. MotionCLIP [68] also sought to align text and action embeddings, advocating for the use of CLIP [59] as a text encoder and incorporating rendered images for additional supervision. This approach showcased the potential to generate actions beyond known distributions and facilitated latent code editing, albeit with limitations in the quality and global translation of the generated action sequences. ACTOR [56] introduced a transformer-based VAE for generating actions from predefined classes in a non-autoregressive fashion. Building on ACTOR’s architecture, TEMOS [57] added a text encoder and produced more diverse action sequences, although it faced challenges with descriptions that fell outside its training distribution. TEACH [7] expanded upon TEMOS [57] by generating sequences of actions from multiple natural language descriptions. The introduction of the large-scale dataset HumanML3D by Guo et al. [25] marked a significant advancement, with the authors also recommending the incorporation of motion length prediction from text to generate motions of appropriate duration. TM2T [26] tackled both text-to-motion and motion-to-text tasks, showing that joint training could yield additional improvements. With the rise of diffusion models, works like MDM [69], MotionDiffuse [80], and MotionLDM [15] have employed diffusion-based approaches for text-to-motion generation, achieving impressive results. However, these models often exhibit lower motion-text matching accuracy and slower inference speeds due to the iterative nature of diffusion processes. Recent studies such as Att-T2M [81] and T2M-GPT [79] have both utilized the foundational VQ-VAE+GPT framework, with the former modeling the spatial relationships of limb parts using a Transformer structure [73] within the VQ-VAE encoder, and the latter leveraging the classic GPT architecture to model the interplay between text and human motion. In contrast to their work, we have explored a more promising RVQ-VAE combined with a novel double-tier GPT structure that offers reduced computational costs and high-quality motion representation. With a straightforward design, this framework achieves superior performance through standard training methodologies, demonstrating its efficacy in the field of text-based human action generation.
@@ 3.2 Motion RVQ-VAE
Inspired by T2M-GPT [79], we also adopt a simple convolutional architecture composed of 1D dilated convolution, residual block [31], and ReLU. Our RVQ-VAE architecture is illustrated in Figure 3. We use 1D convolution with stride 2 and nearest interpolation for temporal downsampling and upsampling respectively. The downsampling rate is thus $l=2^{L}$, where $L$ denotes the number of residual blocks. In our experiments, we set $L=3$ to obtain the downsampling rate $l=8$, each residual block then contains 2 dilated 1D convolution blocks, and the dilated factor is set to 3. More architecture details can be found in the supplementary material.
@@ 3.3 T2M-HiFiGPT
In this section, we propose T2M-HiFiGPT in Figure 2, to autoregressively predict a code stack of motion RVQ-VAE. After we obtain the residual codes extracted by RVQ-VAE. We introduce how our T2M-HiFiGPT efficiently learns the stacked map of discrete codes. We also propose our novel training strategy for T2M-HiFiGPT to prevent the exposure bias [38] problem during the training phase.
@@ Discussion.
In comparison with the vallina GPT model proposed in T2M-GPT [79] and Att-T2M [81], our proposed double-tier GPT structure can theoretically have a better computational advantage given the same number of transformer layers. The computational complexity of vallina GPT is $O(4\times H\times\lfloor T/l\rfloor\times\lfloor T/l\rfloor)$, while our T2M-HiFiGPT is $O(H/2\times\lfloor T/l\rfloor\times\lfloor T/l\rfloor+H/2\times\lfloor T/l\rfloor\times R)$, where $H$ denotes the total transformer layers, which is much less than the vallina GPT model. In the following experimental section, we will compare the detailed parameter counts and inference speed with competitive GPT models [79, 81].
@@ 4.1 Dataset Details
We evaluate our approach using two datasets for text-driven motion generation: HumanML3D [25] and KIT Motion-Language (KIT-ML) [58]. These datasets are widely recognized and utilized within the research community. KIT-ML [58] comprises 3,911 human motion sequences paired with 6,278 textual annotations. The dataset features a vocabulary of 1,623 unique words, excluding variations in capitalization and punctuation. The motion sequences, sourced from the KIT [51] and CMU [16] datasets, are downsampled to 12.5 frames per second (FPS). Each sequence is accompanied by one to four descriptive sentences, with an average description length of about eight words. Following established protocols, we divide the dataset into training, validation, and test sets in an 80%, 5%, and 15% ratio, respectively. The model that performs best on the validation set, as measured by the Fréchet Inception Distance (FID), is then evaluated on the test set. HumanML3D [25], currently the most extensive 3D human motion dataset with textual annotations, contains 14,616 human motions and 44,970 text descriptions. The text corpus encompasses 5,371 distinct words. The motion sequences, derived from AMASS [50] and HumanAct12 [27], undergo specific preprocessing: they are scaled to 20 FPS, motions exceeding 10 seconds are randomly cropped to that duration, re-targeted to a standard human skeletal template, and oriented to initially face the Z+ direction. Each motion is associated with at least three detailed textual descriptions, averaging around 12 words in length. Consistent with prior work [26], we partition this dataset into training, validation, and test sets with the same 80%, 5%, and 15% distribution. Performance on the test set is reported for the model with the best FID score on the validation set.
@@ 4.3 Implementation Details
Table 1: Compare the parameters and computational cost against relevant competitive methods at the reconstruction stage. Time/Frame(ms) denotes the average time cost when generating a motion frame. We measure those metrics by recording the detailed time cost during our testing on the HumanML3D test set.
@@
Table 2: Compare the parameters and computational cost against relevant competitive methods at the generation stage. Time/Frame(ms) denotes the average time cost when generating a motion frame. Time/Token(ms) denotes the average time cost when generating a single discrete motion token by GPT. Both indicators serve as metrics for the response time, and as illustrated in the table, all methods are capable of achieving real-time output at high frame rates.
@@
Table 3: Comparison with the state-of-the-art methods on HumanML3D [25] test set. We compute standard metrics following Guo et al. [25]. For each metric, we repeat the evaluation 20 times and report the average with 95% confidence interval. The bold text indicates the best result.
@@
Table 4: Comparison with the state-of-the-art methods on KIT [58] test set. The experimental settings are the same as Table  LABEL:tab:sota_hm3d. We report the metrics following T2M [25] and repeat 20 times to get the average results with 95% confidence interval. The best results are marked in bold.
@@
Att-T2M [81] $0.413_{\pm.006}$ $0.632_{\pm.006}$ $0.751_{\pm.006}$ $0.870_{\pm.039}$ $3.039_{\pm.021}$ $10.96_{\pm.123}$ $2.281_{\pm.047}$
Ours $\textbf{0.418}_{\pm.005}$ $\textbf{0.637}_{\pm.006}$ $\textbf{0.764}_{\pm.006}$ $\textbf{0.227}_{\pm.013}$ $\textbf{2.828}_{\pm.019}$ $10.71_{\pm.098}$ $1.748_{\pm.036}$

### 24 Dec 2023 v2

@@ Abstract
In this study, we introduce T2M-HiFiGPT, a novel conditional generative framework for synthesizing human motion from textual descriptions. This framework is underpinned by a Residual Vector Quantized Variational AutoEncoder (RVQ-VAE) and a double-tier Generative Pretrained Transformer (GPT) architecture. We demonstrate that our CNN-based RVQ-VAE is capable of producing highly accurate 2D temporal-residual discrete motion representations. Our proposed double-tier GPT structure comprises a temporal GPT and a residual GPT. The temporal GPT efficiently condenses information from previous frames and textual descriptions into a 1D context vector. This vector then serves as a context prompt for the residual GPT, which generates the final residual discrete indices. These indices are subsequently transformed back into motion data by the RVQ-VAE decoder. To mitigate the exposure bias issue, we employ straightforward code corruption techniques for RVQ and a conditional dropout strategy, resulting in enhanced synthesis performance. Remarkably, T2M-HiFiGPT not only simplifies the generative process but also surpasses existing methods in both performance and parameter efficacy, including the latest diffusion-based and GPT-based models. On the HumanML3D [25] and KIT-ML [58] datasets, our framework achieves exceptional results across nearly all primary metrics. We further validate the efficacy of our framework through comprehensive ablation studies on the HumanML3D [25] dataset, examining the contribution of each component. Our findings reveal that RVQ-VAE is more adept at capturing precise 3D human motion with comparable computational demand compared to its VQ-VAE counterparts. As a result, T2M-HiFiGPT enables the generation of human motion with significantly increased accuracy, outperforming recent state-of-the-art approaches such as T2M-GPT [79], Att-T2M [83] and FineMoGen [82].
@@ 1 Introduction
Action generation based on natural language is challenging, mainly because (1) human motion in the real physical world is high-dimensional data with irregular spatial structure, highly dynamic temporal characteristics and many fine details. Direct text-driven synthesis in high-dimensional posture space (such as [45, 71, 69] etc.) is challenging, and slight changes to high-dimensional vectors can lead to visible defects in the generated motion. (2) Actions and text come from different modalities. The model needs to learn the complex mapping relationship from language space to action space across modalities. For this, many works propose using autoencoders [3, 24, 68, 77] and VAE [83, 79, 56, 57, 15] to learn the joint embedding of language and action. MotionClip [68] aligns the action space with the CLIP [59] space. ACTOR [56] and TEMOES [57] respectively proposed Transformer-based VAEs for action-to-action and text-to-action. These works show promising performance under simple descriptions and are limited to generating high-quality actions when the text description becomes long and complex. Guo et al. [25] and TM2T [26] aim to generate action sequences with more challenging text descriptions. However, both methods are not direct, involve three stages of text-to-action generation, and sometimes fail to generate high-quality actions consistent with the text. Recently, diffusion-based models [33] have shown impressive results in image generation [61], then introduced into action generation by MDM [69], MotionDiffuse [80],ReMoDiffuse [81],FineMoGen [82] and MotionLDM [15], and dominate the text-to-action generation task. However, we find that the inference speed improvement of diffusion-based methods [69, 80, 15] may not be as significant as compared to classical methods (such as VQVAE [72]). Recent T2M-GPT [79] and Att-T2M [83] work, both use the VQVAE [72] model as the motion discrete representation; however, we find that real motion capture data contain very rich details, and simply using a single layer of discrete representation is not enough to fully and compactly express these motion details. Simply increasing the number of codebooks will lead to more redundancy, and the motion representation is not compact enough, which will also increase the training difficulty of the generation model due to large codebook sizes and data scarcity. In this work, unlike T2M-GPT [79] and Att-T2M [83], we are inspired by the latest progress in learning discrete representations for generation [5, 19, 20, 22, 49, 61, 72, 75, 78, 18, 74], and studied a simple framework based on Residual Vector Quantization Variational Autoencoder (RVQ-VAE) [78] and Valle-like [74, 40] structure for the compact and highly accurate discrete motion representation and text-to-action generation task. Specifically, we present a two-stage approach for synthesizing actions from textual descriptions. Initially, we employ a standard 1D dilated convolutional network to transform the action sequence into multi-layer residual discrete code indices. Subsequently, drawing inspiration from the speech generation model Valle [74], we introduce an innovative dual-tier GPT model designed to generate motion residual code sequences from pre-trained CLIP [59] text embeddings. Our experimental analysis reveals that the RVQ-VAE model can effectively decrease the downsampling factor of motion data (from 4 to 8), while simultaneously reducing the computational time required for inference. This results in superior reconstruction outcomes on the test set. Moreover, as the sequence length for modeling is halved, the inference duration for the second stage is also proportionately diminished. Despite its straightforward nature, our method is capable of producing high-fidelity motion sequences that are intricately aligned with complex textual descriptions. In practical evaluations, our approach outperformed the GPT-based methods T2M-GPT [79] and Att-T2M [83] as well as recent diffusion-based methods ReMoDiffuse [81] and FineMoGen [82] on the widely recognized datasets HumanML3D [25] and KIT-ML [58]. Notably, on the extensive HumanML3D [25] dataset, our method achieved the highest accuracy, even exceeding the real data in terms of the consistency between text and generated motion, as measured by R-Precision and matching score. Through comprehensive experimentation, we have delved into this domain, and we hope that the insights and conclusions drawn from our work can contribute to future advancements in the field.
@@
    Our findings indicate that the RVQ-VAE model with shared codebook, despite having comparable computational and parameter requirements to the VQ-VAE, exhibits superior performance in reconstructing 3D human motion data, thereby establishing itself as a more effective method for the discrete representation of arbitrary 3D human motion data.
@@ 2 Related Work
Text-based human action generation. Text-based human action generation is a burgeoning field focused on creating 3D human actions from textual descriptions. Early work in this area includes Text2Action [2], which utilized an RNN-based model to produce actions conditioned on brief texts. Language2Pose [3] took this further by employing a curriculum learning strategy to learn a joint embedding space of text and pose, enabling the generation of action sequences from text embeddings. Subsequent research by Ghosh et al. [24] improved upon this by learning separate manifold representations for upper and lower body movements, yielding better results than Language2Pose [3]. MotionCLIP [68] also sought to align text and action embeddings, advocating for the use of CLIP [59] as a text encoder and incorporating rendered images for additional supervision. This approach showcased the potential to generate actions beyond known distributions and facilitated latent code editing, albeit with limitations in the quality and global translation of the generated action sequences. ACTOR [56] introduced a transformer-based VAE for generating actions from predefined classes in a non-autoregressive fashion. Building on ACTOR’s architecture, TEMOS [57] added a text encoder and produced more diverse action sequences, although it faced challenges with descriptions that fell outside its training distribution. TEACH [7] expanded upon TEMOS [57] by generating sequences of actions from multiple natural language descriptions. The introduction of the large-scale dataset HumanML3D by Guo et al. [25] marked a significant advancement, with the authors also recommending the incorporation of motion length prediction from text to generate motions of appropriate duration. TM2T [26] tackled both text-to-motion and motion-to-text tasks, showing that joint training could yield additional improvements. With the rise of diffusion models, works like MDM [69], MotionDiffuse [80], ReMoDiffuse [81],FineMoGen [82] and MotionLDM [15] have employed diffusion-based approaches for text-to-motion generation, achieving impressive results. However, these models often exhibit lower motion-text matching accuracy and slower inference speed due to the iterative nature of diffusion processes. Recent studies such as Att-T2M [83] and T2M-GPT [79] have both utilized the foundational VQ-VAE+GPT framework, with the former modeling the spatial relationships of limb parts using a Transformer structure [73] within the VQ-VAE encoder, and the latter leveraging the classic GPT architecture to model the interplay between text and human motion. In contrast to their work, we have explored a more promising RVQ-VAE combined with a novel double-tier GPT structure that offers reduced computational costs and high-quality motion representation. With a straightforward design, this framework achieves superior performance through standard training methodologies, demonstrating its efficacy in the field of text-based human action generation.
@@ 3.2 Motion RVQ-VAE
Inspired by T2M-GPT [79], we also adopt a simple convolutional architecture composed of 1D dilated convolution, residual block [31], and ReLU. For our RVQ-VAE, we use 1D convolution with stride 2 and nearest interpolation for temporal downsampling and upsampling respectively. The downsampling rate is thus $l=2^{L}$, where $L$ denotes the number of residual blocks. In our experiments, we set $L=3$ to obtain the downsampling rate $l=8$, each residual block then contains 2 dilated 1D convolution blocks, and the dilated factor is set to 3. More architecture details can be found in the supplementary material.
@@ 3.3 T2M-HiFiGPT
In this section, we propose T2M-HiFiGPT in fig. 1, to autoregressively predict a code stack of motion RVQ-VAE. After we obtain the residual codes extracted by RVQ-VAE. We introduce how our T2M-HiFiGPT efficiently learns the stacked map of discrete codes. We also propose our novel training strategy for T2M-HiFiGPT to prevent the exposure bias [62] problem during the training phase.
@@ Discussion.
In comparison with the vallina GPT model proposed in T2M-GPT [79] and Att-T2M [83], our proposed double-tier GPT structure can theoretically have a better computational advantage given the same number of transformer layers. The computational complexity of vallina GPT is $O(4\times H\times\lfloor T/l\rfloor\times\lfloor T/l\rfloor)$, while our T2M-HiFiGPT is $O(H/2\times\lfloor T/l\rfloor\times\lfloor T/l\rfloor+H/2\times\lfloor T/l\rfloor\times R\times R)$, where $H$ denotes the total transformer layers, which is less than the vallina GPT model. In the following experimental section, we will compare the detailed parameter counts and inference speed with competitive GPT models [79, 83].
@@ 4.1 Dataset Details
We evaluate our approach using two datasets for text-driven motion generation: HumanML3D [25] and KIT Motion-Language (KIT-ML) [58]. These datasets are widely recognized and utilized within the research community. KIT-ML [58] comprises 3,911 human motion sequences paired with 6,278 textual annotations. The dataset features a vocabulary of 1,623 unique words, excluding variations in capitalization and punctuation. The motion sequences, sourced from the KIT [51] and CMU [16] datasets, are downsampled to 12.5 frames per second (FPS). Each sequence is accompanied by one to four descriptive sentences, with an average description length of about eight words. Following established protocols, we divide the dataset into training, validation, and test sets in an 80%, 5%, and 15% ratio, respectively. The model that performs best on the validation set, as measured by the Fréchet Inception Distance (FID), is then evaluated on the test set. HumanML3D [25], currently the most extensive 3D human motion dataset with textual annotations, contains 14,616 human motions and 44,970 text descriptions. The text corpus encompasses 5,371 distinct words. The motion sequences, derived from AMASS [50] and HumanAct12 [27], undergo specific preprocessing: they are scaled to 20 FPS, motions exceeding 10 seconds are randomly cropped to that duration, re-targeted to a standard human skeletal template, and oriented to initially face the Z+ direction. Each motion is associated with at least three detailed textual descriptions, averaging around 12 words in length. Consistent with prior work [26], we partition this dataset into training, validation, and test sets with the same 80%, 5%, and 15% distribution. Performance on the test set is reported for the model with the best FID score on the validation set. Note that we use the same motion representation as the prior work [79].
@@ 4.3 Implementation Details
Table 1: Compare the parameters and computational cost against relevant competitive GPT-based methods at the reconstruction stage. Time/Frame(ms) denotes the average time cost when generating a motion frame. We measure those metrics by recording the detailed time cost during our testing on the HumanML3D test set. Our code is implemented using Pytorch.
@@
Table 2: Compare the parameters and computational cost against relevant competitive GPT-based methods at the generation stage. Time/Frame(ms) denotes the average time cost when generating a motion frame. Time/Token(ms) denotes the average time cost when generating a single discrete motion token by GPT. Both indicators serve as metrics for the response time, and as illustrated in the table, all methods are capable of achieving real-time output at high frame rates.
@@
Table 3: Comparison with the state-of-the-art methods on HumanML3D [25] test set. We compute standard metrics following Guo et al. [25]. For each metric, we repeat the evaluation 20 times and report the average with 95% confidence interval. The bold text indicates the best result and the second best is underlined.
@@
ReMoDiffuse [81] $\underline{0.510}_{\pm.002}$ $\underline{0.69}8_{\pm.006}$ $\underline{0.795}_{\pm.004}$ $\underline{0.103}_{\pm.004}$ $\underline{2.974}_{\pm.016}$ $9.018_{\pm.075}$ $1.795_{\pm.043}$
FineMoGen [82] $0.504_{\pm.002}$ $0.690_{\pm.002}$ $0.784_{\pm.002}$ $0.151_{\pm.008}$ $2.998_{\pm.008}$ $9.263_{\pm.094}$ $\underline{2.696}_{\pm.079}$
@@
Table 4: Comparison with the state-of-the-art methods on KIT [58] test set. The experimental settings are the same as table 3. We report the metrics following T2M [25] and repeat 20 times to get the average results with 95% confidence interval. The best results are marked in bold and the second best is underlined.
@@
Att-T2M [83] $0.413_{\pm.006}$ $0.632_{\pm.006}$ $0.751_{\pm.006}$ $0.870_{\pm.039}$ $3.039_{\pm.021}$ $10.96_{\pm.123}$ $\underline{2.281}_{\pm.047}$
ReMoDiffuse [81] $0.427_{\pm.014}$ $0.641_{\pm.004}$ $0.765_{\pm.055}$ $0.155_{\pm.006}$ $2.814_{\pm.012}$ $10.80_{\pm.105}$ $1.239_{\pm.028}$
FineMoGen [82] $0.432_{\pm.006}$ $0.649_{\pm.005}$ $0.772_{\pm.006}$ $\underline{0.178}_{\pm.007}$ $2.869_{\pm.014}$ $10.85_{\pm.115}$ $1.877_{\pm.093}$
Ours $\underline{0.428}_{\pm.006}$ $\textbf{0.650}_{\pm.005}$ $\textbf{0.780}_{\pm.006}$ $0.337_{\pm.012}$ $\textbf{2.777}_{\pm.019}$ $\underline{11.19}_{\pm.099}$ $1.434_{\pm.031}$